OpenStudy (anonymous):

The number of apples required to bake a pie varies directly with the square of the diameter of the pie tin. If 4 apples are required for a pie tin with diameter 8 inches, what is the constant of variation?

OpenStudy (tad1):

n =kd^2 4 = k (64) k = 4/64 k = 1/16